OpenStudy (anonymous):

MEDAL!! What is the slope of the line that passes through (-4,4) and (1,-1)?

pooja195 (pooja195):

\[\huge~\rm~\frac{ x_2-x_1 }{ y_2-y_1 }=slope\]

OpenStudy (anonymous):

y2-y1 = -5 x2-x1= 5

OpenStudy (anonymous):

What is the slope of the line? (0,-3) and (2,3)?

pooja195 (pooja195):

Use the same formula